# Low-depth Clifford circuits approximately solve MaxCut

### v2

@@ Abstract
We introduce a quantum-inspired approximation algorithm for MaxCut based on low-depth Clifford circuits. We start by showing that the solution unitaries found by the adaptive quantum approximation optimization algorithm (ADAPT-QAOA) for the MaxCut problem on weighted fully connected graphs are (almost) Clifford circuits. Motivated by this observation, we devise an approximation algorithm for MaxCut, ADAPT-Clifford, that searches through the Clifford manifold by combining a minimal set of generating elements of the Clifford group. Our algorithm finds an approximate solution of MaxCut on an $N$-vertex graph by building a depth $O(N)$ Clifford circuit. The algorithm has runtime complexity $O(N^{2})$ and $O(N^{3})$ for sparse and dense graphs, respectively, and space complexity $O(N^{2})$, with improved solution quality achieved at the expense of more demanding runtimes. We implement ADAPT-Clifford and characterize its performance on graphs with positive and signed weights. The case of signed weights is illustrated with the paradigmatic Sherrington-Kirkpatrick model, for which our algorithm finds solutions with ground-state mean energy density corresponding to $\sim 94\%$ of the Parisi value in the thermodynamic limit. The case of positive weights is investigated by comparing the cut found by ADAPT-Clifford with the cut found with the Goemans-Williamson (GW) algorithm. For both sparse and dense instances we provide copious evidence that, up to hundreds of nodes, ADAPT-Clifford finds cuts of lower energy than GW. Since good approximate solutions to MaxCut can be efficiently found within the Clifford manifold, we hope our results will motivate to rethink the approach so far used to search for quantum speedup in combinatorial optimization problems.
@@ I Introduction
Quantum computation allows for information processing capacities beyond those attainable in classical digital computing, provided that fault-tolerant quantum computers can be built in practice [1]. While the milestone of break-even error correction of a single logical qubit has been reached [2, 3, 4], the roadmap to scalable fault-tolerant quantum processors extends into the long term. At the same time, the level of noise in near-term quantum devices with modest numbers of qubits precludes their use as general-purpose machines [5, 6].
On the other hand, near-term quantum processors have found a niche in the hybrid quantum-classical model of computation with Variational Quantum Algorithms (VQAs) [7, 8, 9]. These algorithms perform classical optimization of a problem-specific objective function that is evaluated by measuring the output of a parametrized quantum circuit. Through variational search over circuit parameters, a VQA thus seeks a solution circuit that transforms a simple input state in the Hilbert space of problem variables to a superposition of approximate solutions, i.e., configurations that yield near-optimal values of the objective function. In attempts to determine whether they can lead to a speedup over classical algorithms for any useful task, VQAs have been applied to a variety of combinatorial optimization problems. Quantum approximate optimization algorithms (QAOAs) [10, 11] have been in a constant tug of war with classical solvers, with initial indications of putative quantum speedups [12, 13, 14, 15], followed by experimental claims [16] and rebuttals [17], and then further proposals for possible quantum speedup [18]. A byproduct of this large effort has been the definition and construction of quantum-inspired algorithms [19, 20, 21, 22, 23, 24]. These are “dequantized” classical versions of quantum or hybrid quantum-classical algorithms that unveil and exploit previously unrecognized properties or structures in a problem, leading to solution strategies that outperform the best known classical algorithm.
@@ II.2 The quantum approximate optimization algorithm and its adaptive variant
The Quauntum Approximated Optimization Algorithm (QAOA) is a type of variational algorithm [8] that aims to solve combinatorial optimization problems [10]. It is defined by a parametrized quantum circuit with a periodic structure. Each layer of the circuit is given by a product of two unitaries, time evolution under $H_{\rm C}$, followed by time evolution under a mixer Hamiltonian
@@
Not much is known regarding performance guarantees and hardness of approximation for QAOA. The case of constant $p$ has so far been the main focus, as it is the regime of interest for current quantum devices [5]. Ref. [53] gives evidence for a possible quantum advantage for MaxCut on $3$-regular graphs with shallow QAOA. At the same time, it is known that constant $p$ QAOA is bounded away from optimality in sparse graphs [54, 55, 56], as well as in some dense problems where the overlap gap property  [^3] The overlap gap property (OGP) is a property of the geometry of the space of solutions which (very) roughly speaking states that nearly optimal solutions are either close or far apart from each other. The OGP is currently understood as an algorithmic barrier as it implies a large family of algorithms have their best solutions bounded away from optimality. For an extended discussion see Ref. [117]. is known to exist [58]. These results were recently extended to the case of $p\sim\log(N)$ [59]. However these results do not apply directly to the case of $p\sim{\rm poly}(N)$. As a consequence there are no conclusive results on the runtime required for $p\sim{\rm poly}(N)$ QAOA to reach a given approximation ratio, with only loose lower bounds appearing recently [60]. Most studies of QAOA so far have been numerical experiments, meaning that indications of a putative advantage [12] [^4] With further putative evidence coming from numerical experiments relying on techniques to bypass the classical optimization of the variational parameters [110, 102, 18, 118]., have been inconclusive due to the small problem sizes accessible to either quantum implementations or classical simulation.
@@
where the new variational parameter $\gamma_{l}$ is set to a predefined small positive value $\gamma_{0}\sim 0$ [63], and $\mathrm{P}_{\rm OP}$ is an operator pool. The choice of pool is not unique, with different pools being advantageous in different situations [64, 65]. Below we restrict ourselves to the pool
@@ II.3 Clifford circuits and their efficient simulation
Both the QAOA and ADAPT-QAOA ansätze are defined as products of unitaries generated by Pauli strings. When are unitary transformations generated by Pauli strings Clifford unitaries? To answer this question, take $\tilde{P}_{i},\tilde{P}_{j}\in\mathcal{\tilde{P}}$, two distinct Pauli strings that either commute or anticommute by definition. If $W(\theta)=e^{-i\theta\tilde{P}_{j}}$, then $W^{\dagger}\tilde{P}_{i}W=\tilde{P}_{i}$ if $[\tilde{P}_{i},\tilde{P}_{j}]=0$, and $W^{\dagger}\tilde{P}_{i}W=i\tilde{P}_{i}\tilde{P}_{j}$ if $\{\tilde{P}_{i},\tilde{P}_{j}\}=0$ and $\theta=\pm m\frac{\pi}{4}$ with $m\in\mathbb{N}$. We thus see that when a quantum circuit is composed of products of unitaries generated by Pauli strings that do not necessarily commute, it is a Clifford circuit if an only if the parameters of these transformations are integer multiples of ±$\pm\pi/4$. Therefore, if QAOA solution circuits $U(\boldsymbol{\gamma}^{*},\boldsymbol{\beta}^{*})$ are to be Clifford, then the circuit parameters $\boldsymbol{\gamma}^{*}$ and $\boldsymbol{\beta}^{*}$ must be integer multiples of ±$\pm\pi/4$.
@@ II.4 Characterizing the structure of solution unitaries
We analyze the transformation $V$ as an “input-output” channel, with $O$ the input and $O^{\prime}$ the output, and are interested in characterizing the locality, in the Pauli basis, of the output. This can be inferred from the the localization properties of $p_{j}(O;V)$, which we investigate with the Second Renyi entropy, see Ref. [80] and Sec. 2.7 of Ref. [81] [^6] It is interesting to notice the similarities of this quantity with the Stabilizer entropy [119], we restrain ourselves from making any definite claims and leave it for future work.
@@ III Origin of the ADAPT-Clifford algorithm
We consider first the case of graphs with positive weights with $\omega_{i,j}$ from either ${\rm U}[0,1]$, where by ${\rm U}[a,b]$ we denote the uniform distribution in the interval $[a,b]$, or ${\rm Exp}(1)$, the exponential distribution with mean $1$. In Fig. 1a we show the mean approximation ratios for $50$ problem instances with $N=6$ for circuits up to $p=10$ layers. Similar to the observation in Ref. [63], ADAPT-QAOA (green diamonds and circles in Fig. 1a) finds a solution arbitrarily close to the exact solution at sufficiently high but finite $p$, away from the $p\to\infty$ limit where QAOA is guaranteed to reach the exact solution. In Fig. 1b we show the expectation value over instances of $\overline{\mathcal{S}}$, $\mathbb{E}[\overline{\mathcal{S}}]$. The ADAPT-QAOA solution circuits that lead to $\alpha\to 1$ in Fig. 1a display $\mathbb{E}[\overline{\mathcal{S}}]\to 0$, indicating they might be Clifford circuits. In contrast, the QAOA solution unitaries show $\mathbb{E}[\overline{\mathcal{S}}]>0$ with a tendency towards the typical value, $\log(4^{-N})$, with increasing depth, in agreement with previous works using other indicators [85, 86].
@@ IV.1 Details of the algorithm
We design ADAPT-Clifford so as to exploit the observations summarized at the end of Sec. III in preparing a stabilizer state $|\Psi\rangle$ satisfying the properties described above. In particular, ADAPT-Clifford prepares the stabilizer state
$|\Psi\rangle=\left[\prod_{r=2}^{N-1}e^{i\frac{\pi}{4}Z_{a^{(r)}}Y_{b^{(r)}}}\right]e^{i\frac{\pi}{4}Z_{b^{(1)}}Y_{a^{(1)}}}Z_{k}\mathrm{H}^{\otimes N}|0\rangle^{\otimes N},$ (16)
where $k$ is the position of a qubit chosen arbitrarily, $a^{(r)}\in\mathbf{a}^{(r)}$ and $b^{(r)}\in\mathbf{b}^{(r)}$ are indices denoting the positions of “active” and “inactive” qubits, respectively, and $\mathbf{a}^{(r)}$ and $\mathbf{b}^{(r)}$ are vectors storing the positions of all the active and inactive qubits at step $r$. We call a qubit active if a Pauli gate has been applied to it, otherwise it is inactive.
@@
    After all $N$ steps are completed, we perform a measurement in the computational basis. From the output bit string, $\mathbf{z}_{\rm out}$, we readout the approximate maximal cut of the graph as $(\mathcal{A},\overline{\mathcal{A}})$ with $\mathcal{A}=\{z_{i}\in\mathbf{z}_{\rm out}|z_{i}=0,i=1,..,N\}$ and $\mathcal{A}=\{z_{i}\in\mathbf{z}_{\rm out}|z_{i}=1,i=1,..,N\}$.
@@
where the $\mathrm{S}_{l}$, $\mathrm{H}_{l}$, are the phase and Hadamard gates acting on the $l$-th qubit, ${\rm CNOT}_{l,m}$ is the controlled NOT gate, with qubit $l$ and qubit $m$ as control and target qubits, respectively. Furthermore one can write $R_{x}^{(l)}(-\pi/2)=\mathrm{H}_{l}^{YZ}Z_{l}$ with $\mathrm{H}_{l}^{YZ}$ a variant of the Hadamard gate which swaps the $y$- and $z$-axes. For the interested reader, we work through the operations of our algorithm for two small examples in App. D.
@@ V Algorithm performance on weighted complete graphs
We have implemented the ADAPT-Clifford algorithm using the fast stabilizer circuit simulator Stim [29]. Our implementation is available at  [^7] https://github.com/manuelmz/adapt-clifford. Although we have chosen this simulator to implement our algorithm, any stabilizer circuit simulator which supports interactivity, that is, where expectation values of Pauli strings can be computed and the circuit modified according to the results, could be used to implement the algorithm.
@@ V.1 The case of positive weights
We begin studying the performance of the randomized approach. We draw a parallel between the random initialization of ADAPT-Clifford and the rounding step of GW, and thus assess the performance of the randomized ADAPT-Clifford by direct comparison with GW. We solved $100$ different problem instances for graph sizes up to $N=1000$ with both algorithms. In Fig. 2a we show the normalized mean minimum energy, $\mathbb{E}[E_{\rm min}]/N$, of the solutions obtained with randomized ADAPT-Clifford (green circles) and the ones obtained with GW (light grey circles). Notice that our randomized ADAPT-Clifford almost always produces a solution of lower energy expectation than GW. These observations can be further verified with the mean difference of the minimum energy found, $\mathbb{E}[E_{\rm min}^{\rm Cliff}-E_{\rm min}^{\rm GW}]$, which we show in Fig. 2b. Since our randomized ADAPT-Clifford consistently beats GW, we expect it to have a performance guarantee for typical instances of positively weighted complete graphs above that of GW for the general problem. We discuss the methodology used to estimate it in App. E. We find $\overline{\alpha}^{\rm r}\approx 0.8986$ a value which confirms our intuition and sets a lower bound for the expected performance of the deterministic ADAPT-Clifford.
@@
To further verify this observation, the inset of Fig. 3c shows the mean energy difference, $\mathbb{E}[E_{\rm min}^{\rm Cliff}-E_{\rm min}^{\rm GW}]$, between the solution found with our algorithm and the one found with GW, with the magenta dotted line indicating $\mathbb{E}[E_{\rm min}^{\rm Cliff}-E_{\rm min}^{\rm GW}]=0$, that is, equal quality cuts on average. It is seen that ADAPT-Clifford performs increasingly better than GW with fixed $\mathcal{I}$ as problem size is increased. To quantify the approximation quality of ADAPT-Clifford, we estimate the average approximation ratio of the deterministic ADAPT-Clifford on this family of graphs to be $\overline{\alpha}=0.9686$ – see App. E for details.
While Fig. 3c shows that $\mathcal{I}=10^{5}$ rounding steps are needed for the GW algorithm to match the approximation quality of the deterministic ADAPT-Clifford for problem sizes up to $N=200$, the data in the inset imply that $\mathcal{I}$ may in fact need to scale with $N$ for the GW algorithm to compete with ADAPT-Clifford. Therefore, although the expected runtime of standard GW $O(N^{3.5})$, see [30, 91] and references therein, compared to $O(N^{4})$ for ADAPT-Clifford on complete graphs, our benchmarks are inconclusive as to which of the two algorithms is faster.

### v3

@@ Abstract
We introduce a quantum-inspired approximation algorithm for MaxCut based on low-depth Clifford circuits. We start by showing that the solution unitaries found by the adaptive quantum approximation optimization algorithm (ADAPT-QAOA) for the MaxCut problem on weighted fully connected graphs are (almost) Clifford circuits. Motivated by this observation, we devise an approximation algorithm for MaxCut, ADAPT-Clifford, that searches through the Clifford manifold by combining a minimal set of generating elements of the Clifford group. Our algorithm finds an approximate solution of MaxCut on an $N$-vertex graph by building a depth $O(N)$ Clifford circuit. The algorithm has runtime complexity $O(N^{2})$ and $O(N^{3})$ for sparse and dense graphs, respectively, and space complexity $O(N^{2})$, with improved solution quality achieved at the expense of more demanding runtimes. We implement ADAPT-Clifford and characterize its performance on graphs with positive and signed weights. The case of signed weights is illustrated with the paradigmatic Sherrington-Kirkpatrick model, for which our algorithm finds solutions with ground-state mean energy density corresponding to $\sim 94\%$ of the Parisi value in the thermodynamic limit. The case of positive weights is investigated by comparing the cut found by ADAPT-Clifford with the cut found with the Goemans-Williamson (GW) algorithm. For both sparse and dense instances we provide copious evidence that, up to hundreds of nodes, ADAPT-Clifford finds cuts of lower energy than GW.
@@ I Introduction
Near-term quantum processors have found a niche in the hybrid quantum-classical model of computation with Variational Quantum Algorithms (VQAs) [1, 2, 3]. These algorithms perform classical optimization of a problem-specific objective function that is evaluated by measuring the output of a parametrized quantum circuit. Through variational search over circuit parameters, a VQA thus seeks a solution circuit that transforms a simple input state in the Hilbert space of problem variables to a superposition of approximate solutions, i.e., configurations that yield near-optimal values of the objective function. In attempts to determine whether they can lead to a speedup over classical algorithms for any useful task, VQAs have been applied to a variety of combinatorial optimization problems.
Quantum approximate optimization algorithms (QAOAs) [4, 5] have been in a constant tug of war with classical solvers, with initial indications of putative quantum speedups [6, 7, 8, 9], followed by experimental claims [10] and rebuttals [11], and then further proposals for possible quantum speedup [12]. A byproduct of this large effort has been the definition and construction of quantum-inspired algorithms [13, 14, 15, 16, 17, 18]. These are “dequantized” classical versions of quantum or hybrid quantum-classical algorithms that unveil and exploit previously unrecognized properties or structures in a problem, leading to solution strategies that outperform the best known classical algorithm.
@@ II.2 The quantum approximate optimization algorithm and its adaptive variant
The Quantum Approximate Optimization Algorithm (QAOA) is a type of variational algorithm [2] that aims to solve combinatorial optimization problems [4]. It is defined by a parametrized quantum circuit with a periodic structure. Each layer of the circuit is given by a product of two unitaries, time evolution under $H_{\rm C}$, followed by time evolution under a mixer Hamiltonian
@@
Not much is known regarding performance guarantees and hardness of approximation for QAOA. The case of constant $p$ has so far been the main focus, as it is the regime of interest for current quantum devices [47]. Ref. [48] gives evidence for a possible quantum advantage for MaxCut on $3$-regular graphs with shallow QAOA. Refs. [49, 50] provide evidence that both $p=1$ and large-depth QAOA output states with bit string probabilities following Boltzman distributions, rendering sampling classically hard. At the same time, it is known that constant $p$ QAOA is bounded away from optimality in sparse graphs [51, 52, 53], as well as in some dense problems where the overlap gap property  [^3] The overlap gap property (OGP) is a property of the geometry of the space of solutions which (very) roughly speaking states that nearly optimal solutions are either close or far apart from each other. The OGP is currently understood as an algorithmic barrier as it implies a large family of algorithms have their best solutions bounded away from optimality. For an extended discussion see Ref. [117]. is known to exist [55]. These results were recently extended to the case of $p\sim\log(N)$ [56]. However these results do not apply directly to the case of $p\sim{\rm poly}(N)$. As a consequence there are no conclusive results on the runtime required for $p\sim{\rm poly}(N)$ QAOA to reach a given approximation ratio, with only loose lower bounds appearing recently [57]. Most studies of QAOA so far have been numerical experiments on different families of problem instances, two examples are Erdos-Renyi graphs [6] and $3$- and $4$-regular graphs (weighted and unweighted) [58]. Importantly, any indication of a putative advantage in this type of studies, has been inconclusive due to the small problem sizes accessible to either quantum implementations or classical simulation  [^4] With further putative evidence coming from numerical experiments relying on techniques to bypass the classical optimization of the variational parameters [58, 102, 12, 118] or to steer the optimization process in a more efficient and effective manner [119]..
@@
where the new variational parameter $\gamma_{l}$ is set to a predefined small positive value $\gamma_{0}\sim 0$ [61], $\mathrm{P}_{\rm OP}$ is an operator pool, and $|\psi_{l-1}\rangle$ is the state resulting from the application of the ADAPT-QAOA solution circuit with only $l-1$ layers. The choice of pool is not unique, with different pools being advantageous in different situations [62, 63]. Below we restrict ourselves to the pool
@@ II.3 Clifford circuits and their efficient simulation
Both the QAOA and ADAPT-QAOA ansätze are defined as products of unitaries generated by Pauli strings. When are unitary transformations generated by Pauli strings Clifford unitaries? To answer this question, take $\tilde{P}_{i},\tilde{P}_{j}\in\mathcal{\tilde{P}}$, two distinct Pauli strings that either commute or anticommute by definition. Further consider the unitary $W(\theta)=e^{-i\theta\tilde{P}_{j}}$, then $W^{\dagger}\tilde{P}_{i}W=\tilde{P}_{i}$ if $[\tilde{P}_{i},\tilde{P}_{j}]=0$, and $W^{\dagger}\tilde{P}_{i}W=i\tilde{P}_{i}\tilde{P}_{j}$ if $\{\tilde{P}_{i},\tilde{P}_{j}\}=0$ and $\theta=\pm m\frac{\pi}{4}$ with $m\in\mathbb{N}$. We thus see that when a quantum circuit is composed of products of unitaries generated by Pauli strings that do not necessarily commute, it is a Clifford circuit if an only if the parameters of these transformations are integer multiples of ±$\pm\pi/4$. Therefore, if QAOA solution circuits $U(\boldsymbol{\gamma}^{*},\boldsymbol{\beta}^{*})$ are to be Clifford, then the circuit parameters $\boldsymbol{\gamma}^{*}$ and $\boldsymbol{\beta}^{*}$ must be integer multiples of ±$\pm\pi/4$.
@@ II.4 Characterizing the structure of solution unitaries
We analyze the transformation $V$ as an “input-output” channel, with $O$ the input and $O^{\prime}$ the output, and are interested in characterizing the locality, in the Pauli basis, of the output. This can be inferred from the the localization properties of $p_{j}(O;V)$, which we investigate with the Second Renyi entropy, see Ref. [79] and Sec. 2.7 of Ref. [80]
@@
Eq. (12) has a resemblance to the stabilizer Renyi entropy [81]. Although the latter quantifies the nonstabilizerness of a multi-qubit state, the expression in Eq. (12) directly looks at nonCliffordness of the transformation. As such, one might interpret it as the operator space counterpart to the stabilizer Renyi entropy, and we expect both quantities to have similar behaviors, that is, if for a multi-qubit state $V|\psi\rangle^{\otimes N}$ the stabilizer Renyi entropy is high/low, then Eq. (12) for some input Pauli string $O$ and the same unitary $V$ will be high/low.
@@ III Origin of the ADAPT-Clifford algorithm
We consider first the case of graphs with positive weights with $\omega_{i,j}$ from either ${\rm U}[0,1]$, where by ${\rm U}[a,b]$ we denote the uniform distribution in the interval $[a,b]$, or ${\rm Exp}(1)$, the exponential distribution with mean $1$. In Fig. 1a we show the mean approximation ratios for $50$ problem instances with $N=6$ for circuits up to $p=10$ layers. Similar to the observation in Ref. [61], ADAPT-QAOA (green diamonds and circles in Fig. 1a) finds a solution arbitrarily close to the exact solution at sufficiently high but finite $p$, away from the $p\to\infty$ limit where QAOA is guaranteed to reach the exact solution. In Fig. 1b we show the expectation value over instances of $\overline{\mathcal{S}}$, $\mathbb{E}[\overline{\mathcal{S}}]$. The ADAPT-QAOA solution circuits that lead to $\alpha\to 1$ in Fig. 1a display $\mathbb{E}[\overline{\mathcal{S}}]\to 0$, indicating they might be Clifford circuits. In contrast, the QAOA solution unitaries show $\mathbb{E}[\overline{\mathcal{S}}]>0$ with a tendency towards the typical value, $\log(4^{-N})$, with increasing depth, in agreement with previous works using other indicators [84, 85, 86].
@@ IV.1 Details of the algorithm
We design ADAPT-Clifford so as to exploit the observations summarized at the end of Sec. III to prepare a stabilizer state $|\Psi\rangle$ with the general form given in Eq. (15). This is done in a greedy manner, where after a choice of an initial seed, at every step the best local update is performed. As such, at an intermediate step $0<r\leq N$ we label qubits as active and inactive. A qubit is active if a Pauli gate has been applied to it, otherwise it is inactive, and $a^{(r)}\in\mathbf{a}^{(r)}$ and $b^{(r)}\in\mathbf{b}^{(r)}$ are indices denoting the positions of “active” and “inactive” qubits, respectively, and $\mathbf{a}^{(r)}$ and $\mathbf{b}^{(r)}$ are vectors storing the positions of all the active and inactive qubits at step $r$.
@@
    After all $N$ steps are completed, we perform a measurement in the computational basis. From the output bit string, $\mathbf{z}_{\rm out}$, we readout the approximate maximal cut of the graph as $(\mathcal{A},\overline{\mathcal{A}})$ with $\mathcal{A}=\{z_{i}\in\mathbf{z}_{\rm out}|z_{i}=0,i=1,..,N\}$ and $\overline{\mathcal{A}}=\{z_{i}\in\mathbf{z}_{\rm out}|z_{i}=1,i=1,..,N\}$.
After the above $N$ steps are completed, the resulting stabilizer state $|\Psi\rangle$ encoding the solution has the form
$|\Psi\rangle=\left[\prod_{r=2}^{N-1}e^{i\frac{\pi}{4}Z_{\tilde{l}}Y_{b^{(r)}}}\right]e^{i\frac{\pi}{4}Z_{j}Y_{k}}Z_{k}\mathrm{H}^{\otimes N}|0\rangle^{\otimes N}.$ (20)
@@
where the $\mathrm{S}_{l}$, $\mathrm{H}_{l}$, are the phase and Hadamard gates acting on the $l$-th qubit, ${\rm CNOT}_{l,m}$ is the controlled NOT gate, with qubit $l$ and qubit $m$ as control and target qubits, respectively. Furthermore one can write $R_{x}^{(l)}(-\pi/2)=\mathrm{H}_{l}^{YZ}Z_{l}$ with $\mathrm{H}_{l}^{YZ}$ a variant of the Hadamard gate which swaps the $y$- and $z$-axes. For the interested reader, we work through the operations of our algorithm for two small examples in App. F.
@@ V Algorithm performance on weighted complete graphs
We have implemented the ADAPT-Clifford algorithm using the fast stabilizer circuit simulator Stim [23]. Our implementation is available at [90]. Although we have chosen this simulator to implement our algorithm, any stabilizer circuit simulator which supports interactivity, that is, where expectation values of Pauli strings can be computed and the circuit modified according to the results, could be used to implement the algorithm.
@@ V.1 The case of positive weights
We begin studying the performance of the randomized approach. We draw a parallel between the random initialization of ADAPT-Clifford and the rounding step of GW, and thus assess the performance of the randomized ADAPT-Clifford by direct comparison with GW. We solved $100$ different problem instances for graph sizes up to $N=1000$ with both algorithms. In Fig. 2a we show the normalized mean minimum energy, $\mathbb{E}[E_{\rm min}]/N$, of the solutions obtained with randomized ADAPT-Clifford (green circles) and the ones obtained with GW (light grey circles). Notice that our randomized ADAPT-Clifford almost always produces a solution of lower energy expectation than GW. These observations can be further verified with the mean difference of the minimum energy found, $\mathbb{E}[E_{\rm min}^{\rm Cliff}-E_{\rm min}^{\rm GW}]$, which we show in Fig. 2b. Since our randomized ADAPT-Clifford consistently beats GW, we expect it to have a performance guarantee for typical instances of positively weighted complete graphs above that of GW for the general problem. We discuss the methodology used to estimate it in App. G. We find $\overline{\alpha}^{\rm r}\approx 0.8986$ a value which confirms our intuition and sets a lower bound for the expected performance of the deterministic ADAPT-Clifford.
@@
To further verify this observation, the inset of Fig. 3c shows the mean energy difference, $\mathbb{E}[E_{\rm min}^{\rm Cliff}-E_{\rm min}^{\rm GW}]$, between the solution found with our algorithm and the one found with GW, with the magenta dotted line indicating $\mathbb{E}[E_{\rm min}^{\rm Cliff}-E_{\rm min}^{\rm GW}]=0$, that is, equal quality cuts on average. It is seen that ADAPT-Clifford performs increasingly better than GW with fixed $\mathcal{I}$ as problem size is increased. To quantify the approximation quality of ADAPT-Clifford, we estimate the average approximation ratio of the deterministic ADAPT-Clifford on this family of graphs to be $\overline{\alpha}=0.9686$ – see App. G for details.
To complement our benchmarks we performed a time to solution (TTS) study of GW with variable $\mathcal{I}$, randomized ADAPT-Clifford and deterministic ADAPT-Clifford, the details are shown in App. D. While Fig. 3c shows that $\mathcal{I}=10^{5}$ rounding steps are needed for the GW algorithm to match the approximation quality of the deterministic ADAPT-Clifford for problem sizes up to $N=200$, the data in the inset imply that $\mathcal{I}$ may in fact need to scale with $N$ for the GW algorithm to compete with ADAPT-Clifford. However, in App. D we show that the TTS of GW remains constant with $I$ up to $I=10^{3}$ and only slightly increased for larger values of $\mathcal{I}$, and even at $\mathcal{I}=10^{5}$ the TTS of GW is faster than deterministic ADAPT-Clifford. In fact, we empirically verify the $O(N^{3})$ runtime scaling of GW (see App. D and [24, 91] and references therein).
However, we also find an empiric runtime scaling for randomized ADAPT-Clifford of $O(N^{2.7})$, indicating an advantage. The comparison here is more involved, as GW with $\mathcal{I}>10^{1}$ already produces a better solution than randomized ADAPT-Clifford. However the Cholesky decomposition performed as part of GW is usually executed as a multi-core operation in most numerical linear algebra libraries. On the contrary, our implementation of ADAPT-Clifford runs on a single core, thus one could then easily improve the quality of solution without sacrificing the TTS or runtime scaling by executing the algorithm for different initial position in parallel. We thus believe, this variant of the algorithm does offer an advantage over GW.